The perimeter of a rectangle was 120. If the length is four times the width, what is the true length measurement.
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 60
4w + w = 60
5w = 60
w = 12 and L = 48
And...checking
L + w = 60